OpenStudy (egracer):

Given the functions a(x) = 3x − 12 and b(x) = x − 9, solve a[b(x)] and select the correct answer below.

OpenStudy (egracer):

the choices are a[b(x)] = 3x2 − 21 a[b(x)] = 3x2 − 39 a[b(x)] = 3x − 21 a[b(x)] = 3x − 39

Vocaloid (vocaloid):



Vocaloid (vocaloid):

basically, you want to take b(x) and substitute it for x in the expression a(x) = 3(x) - 12 a(x) = 3x-12 a(b(x)) = 3*(b(x)) - 12 = 3(x-9) - 12